OpenStudy (anonymous):

Let f(x) = 12 over 4 x + 2 Find f(-1). -6 -2 2 6

OpenStudy (anonymous):

\[f(x)=\frac{12}{4x+2}\] \[f(-1)=\frac{12}{4\times (-1)+2}\] is a start

OpenStudy (anonymous):

so then it would be 12 over -2?

OpenStudy (anonymous):

oh so -6! Thanks!

OpenStudy (anonymous):

yes, that is it

OpenStudy (anonymous):

yw